A=2I+3JAND B=5I-6J THE MAGNITUDE OF A+B IS                           a)4 Units B)10 units C)root 58 units D) root 61 units
 

Grade:11

1 Answers

Tanishq Gupta
2 years ago
Hi Vatsal,
A + B = 7I-3J
And Magnitude of A+B=root (A2+B2)
Therefore correct answer is root 58 units (option C)
Thank You.